$SPAD/src/input drawcfn.input

The Axiom Team December 3, 2016

Abstract

Contents

2

— * —

-- Two dimensional function plots
f(x:SF):SF == sin(11*x)
draw(f,0..2*%pi)
draw(f,0..2*%pi,adaptive == false,title == "Non-adaptive plot")
draw(f,0..2*%pi,toScale == true,title == "Drawn to scale")
draw(f,0..2*%pi,coordinates == polar,title == "Polar plot")
g(x:SF):SF == tan(x)
draw(g,-6..6,title == "Clipping on")
draw(g,-6..6,clip == false,title == "Clipping off")
-- Parametric plane curves
f1(t:SF):SF == 9*sin(3*t/4)
f2(t:SF):SF == 8*sin(t)
draw(curve(f1,f2),-4*%pi..4*%pi,toScale == true,title == "Lissajous curve")
g1(t:SF):SF == sin(5*t)
g2(t:SF):SF == t
draw(curve(g1,g2),0..2*%pi,title == "Parametric curve")
draw(curve(g1,g2),0..2*%pi,_
           coordinates == polar,title == "Parametric polar curve")
h1(t:SF):SF == t
h2(t:SF):SF == 2
draw(curve(h1,h2),-3..3,coordinates == parabolic,title == "Parabolic plot")
-- Parametric space curves
i1(t:SF):SF == sin(t)*cos(3*t/5)
i2(t:SF):SF == cos(t)*cos(3*t/5)
i3(t:SF):SF == cos(t)*sin(3*t/5)
draw(curve(i1,i2,i3),0..15*%pi,title == "Parametric curve")
draw(curve(i1,i2,i3),0..15*%pi,tubeRadius == .15,title == "Tube around curve")
j1(t:SF):SF == 4
j2(t:SF):SF == t
draw(curve(j1,j2,j2),-9..9,coordinates == cylindrical,title == "Spiral")
k1(t:SF):SF == 1
k2(t:SF):SF == t/7
k3(t:SF):SF == t/5
draw(curve(k1,k2,k3),-70*%pi..70*%pi,coordinates == spherical,_
           title == "Gnarly curve on sphere")

-- Three dimensional function plots
l(x:SF,y:SF):SF == cos(x*y)
draw(l,-3..3,-3..3)
colorFunction1(x:SF,y:SF):SF == x
draw(l,-3..3,-3..3,colorFunction == colorFunction1)
colorFunction2(x:SF,y:SF):SF == x**2 + y**2
draw(l,-3..3,-3..3,colorFunction == colorFunction2)
colorFunction3(x:SF,y:SF,z:SF):SF == -z
draw(l,-3..3,-3..3,colorFunction == colorFunction3)
m(u:SF,v:SF):SF == 1
draw(m,0..2*%pi,0..%pi,coordinates == spherical,title == "Sphere")
draw(m,0..2*%pi,0..%pi,coordinates == spherical,_
           colorFunction == colorFunction2,title == "Sphere")
draw(m,-%pi/2..%pi/2,0..2*%pi,coordinates == oblateSpheroidal(1$SF),_
           title == "Oblate spheroidal coordinates")
-- Parametric surfaces
n1(u:SF,v:SF):SF == 5*sin(u)*cos(v)
n2(u:SF,v:SF):SF == 4*sin(u)*sin(v)
n3(u:SF,v:SF):SF == 3*cos(u)
draw(surface(n1,n2,n3),0..%pi,0..2*%pi,title == "Which came first?")

———-

References^{[1] nothing}

5